OpenStudy (anonymous):

2. Simplify 1/4(24-16x) (1 point)

OpenStudy (mathstudent55):

Do you mean this? \( \dfrac{1}{4}(24 - 16x) \) If so, distribute the fraction 1/4 by the terms in the parentheses.

OpenStudy (anonymous):

6-4x

OpenStudy (mathstudent55):

Here's a similar example. Follow it and then try yours. \( \dfrac{1}{2}(8 - 16x) \) \( = \dfrac{1}{2}(8) - \dfrac{1}{2}(16x) \) \( = 4 - 8x\)

OpenStudy (mathstudent55):

@ksrakhilesh Giving out answers is against the code of conduct. In addition, you're not helping the person who asked the question understand the problem or learn anything from the problem.

OpenStudy (anonymous):

ok iam sorry